A rectangular lawn has a length that is 3 yards greater than the width. The area of the lawn is 130 square yards.
Note: The formula for the area of a rectangle is Area = length \times width .
a) Write the polynomial equation for the area of the lawn. Use the variable x to represent the width of the lawn.

Answer by mananth(16946)

let the width of the lawn be x
length =x+3
x*(x+3)=130
x^2+3x-130=0
x^2+13x-10x-130=0
x(x+13)-10(x+13)=0
(x+13)(x-10)=0
so x= 10